Determine the pH (to two decimal places) of the following solutions. A table of pKa values can be found here. If the 5% approximation is valid, use the assumption to compute pH.

6.80×10-3 M butanoic acid

Tries 0/13



9.74×10-1 M formic acid

Tries 0/13



8.47×10-3 M methylamine

Tries 0/13



6.59×10-3 M triethylamine
Answer #1

ic Acid PKa = 4.32 6-80 s .41.08 3.49 the out om 3.49 PKa - 3.75 a2 815 572 xL = 1,830千 (3.36 +2.07 =(4- 14-27:5 I285pH = 14 4- á (s.as -sx)) (825+21e) 14ーマ = 14-dx 543 11.485